OpenStudy (anonymous):

help please Solve each of these systems using the substitution method. Make sure to check your answers ! 2x + 3y = -4 x − 2y = 5 Answer a. x = 1, y = -2 b. x = -1, y = 2 c. x = 1, y = 2 d. x = -2, y = -1 question 2 2x − y = -2 3y = 6 − 5x Answer a. x = 0, y = 2 b. x = 2, y = 0 c. x = 2, y = 2 d. x = 0, y = -2 question 3 2x − y = 4 6x − 3y = 7 Answer a. y = 2x - 4 b. No solution c. x = 2, y = -4 d. x = 2, y = 5/3 question 4 2x − y = 3 4x = 6 + 2y Answer a. Infinitely many solutions b. x = 2, y = 1 c.

OpenStudy (anonymous):

1 is a

OpenStudy (anonymous):

2 is a

OpenStudy (anonymous):

3 is b

OpenStudy (anonymous):

4 is b

OpenStudy (texaschic101):

1. x - 2y = 5 x = 2y + 5 now sub this into the other equation 2x + 3y = - 4 2(2y + 5) + 3y = - 4 4y + 10 + 3y = - 4 7y = - 4 - 10 7y = - 14 y = -14/7 y = - 2 now sub this into either of the equations x - 2y = 5 x - 2(-2) = 5 x + 4 = 5 x = 5 - 4 x = 1 ANSWER : (1,-2) I just did one to show you how to do these kinds of problems. You first have to get a variable by itself. Then sub that answer into the other equation. Once that variable is found, sub that answer into either equation to find the other variable. You can then check your answer by subbing in known variables back into either of the equations and if they equal, it is correct. If they don't equal, then a mistake made.